# IEEEexample:BSTcontrol

### v1

@@ 2.1 Federated Learning
Federated learning (FL) enables multiple sites to jointly train a global model without sharing data, thereby alleviating data privacy concerns [8]. Although FL has witnessed success in medical and surgical vision tasks via federated averaging [9], balanced weights sharing strategies [10], dynamic weight averaging [21], style transfer [22], and contrastive learning [11], these works focus on obtaining a single global model, which struggles to perform well across sites. This is prominent in surgical scenarios where different clinical settings, procedures, imaging systems, and patients introduce high heterogeneity in data distribution. In light of this limitation, personalized federated learning (PFL) is emerging as an appealing alternative. Unlike FL, which creates a single global model, PFL aims to produce individualized models that cater to the local distributions of each site’s data to enhance performance with privacy protection [12]. Many PFL approaches decouple the model architecture into global and personalized layers, where the global parameters are aggregated and shared via FedAvg [8] at the server, while the personalized ones remain local. The global layer can be composed of prediction head layers [13], batch normalization layers [14], or parts of channels in convolutional layers [15]. Some methods explore and utilize knowledge among different sites. For instance, FedDP [16] employs the inconsistencies in ground-truth annotations from various sites to calibrate local training. The pFedLA [17] utilizes hypernetworks at the server, instead of distance metrics, to learn the similarities for all parameters between different sites. Unlike these methods, our novel approach focuses on separating global and personalized parameters following a multi-headed perspective. Moreover, it introduces shape similarity and appearance discrepancy priors in SIS for global and personalized parameters, respectively, thus providing site-independent shape and site-specific appearance information for each site.
@@ 5 Conclusions
This paper presents PFedSIS, a novel PFL method for SIS, capitalizing on visual trait priors of appearance disparity and instrument shape similarity. It first decouples the embedding heads of MSA and channels of convolution layers into personalized and global via GPD. Then, APE enhances personalized parameters tailored to each site’s appearance representation and utilizes inter-site appearance inconsistency via hypernetwork for personalized parameters update, while SGE enables global parameters to maintain shape information via cross-style consistency and effectively share mutual shape information via shape-similarity update. Experimental results on three publicly available datasets show that the proposed PFedSIS achieves superior performance in both quantitative assessment and visual interpretation, while its computation requirements during prediction remain low and comparable to other methods. Future work will involve collaborating with other surgical disciplines, such as ophthalmic surgery and neurosurgery, to develop personalized foundational models for surgical instrument segmentation.

### v2

@@ 2.1 Federated Learning
Federated learning (FL) enables multiple sites to jointly train a global model without sharing data, thereby alleviating data privacy concerns [8]. Although FL has witnessed success in medical and surgical vision tasks via federated averaging [9], balanced weights sharing strategies [10], dynamic weight averaging [21], style transfer [22], and contrastive learning [11], these works focus on obtaining a single global model, which struggles to perform well across sites. This is prominent in surgical scenarios where different clinical settings, procedures, imaging systems, and patients introduce high heterogeneity in data distribution. In light of this limitation, personalized federated learning (PFL) is emerging as an appealing alternative. Unlike FL, which creates a single global model, PFL aims to produce individualized models that cater to the local distributions of each site’s data to enhance performance with privacy protection [12]. Many PFL approaches decouple the model architecture into global and personalized layers, where the global parameters are aggregated and shared via FedAvg [8] at the server, while the personalized ones remain local. The global layer can be composed of prediction head layers [13], batch normalization layers [14], or parts of channels in convolutional layers [15]. Some methods explore and utilize knowledge among different sites. For instance, FedDP [16] employs the inconsistencies in ground-truth annotations from various sites to calibrate local training. The pFedLA [17] utilizes hypernetworks at the server, instead of distance metrics, to learn the similarities for all parameters between different sites. Unlike these methods, our novel approach focuses on separating global and personalized parameters following a multi-headed perspective. Moreover, it introduces shape similarity and appearance discrepancy priors for global and personalized parameters, respectively, thus providing site-independent shape and site-specific appearance information for each site.
@@ 4.4.2 Impact of Different Personalized Settings in GPD
5 Discussion
Surgical instrument segmentation is essential for robotics-assisted surgery and computer-aided interventions. It can enhance surgical precision, reduce the risk of accidental damage to vital structures and elevate understanding of operating field. Developing an expert segmentation model necessitates a substantial amount of annotated data and relying exclusively on data from a single clinical site is inadequate to meet training requirements, while the bias inherent in single-centre data impairs the model’s generalization ability. Moreover, ethical and privacy policies must be satisfied with controlled data access.
This paper represents the initial effort to use PFL for surgical instrument segmentation, which enables collaborative training without compromising patient privacy while ensuring site personalization through the aggregation of knowledge from various sites. Unlike other PFL approaches, our PFedSIS proposes a GPD to consider head-wise personalization for the multi-headed self-attention mechanism, allowing for more granular customization of segmentation models. Furthermore, PFedSIS propose APE and SGE based on appearance discrepancy and shape similarity priors to exploit inter-site inconsistency and consistency, respectively, achieving efficient segmentation performance on each site. On the one hand, APE utilizes appearance regulation to guide personalized parameters toward the local site’s appearance distribution and build individual hypernetworks for each site to customize personalized parameter updates. On the other hand, SGE applies cross-style shape consistency to the input and its segmentation map, which enables global parameters to focus on mutual shape information across different sites, and preserves global parameters crucial for shape information via shape-similarity update. Comparison experimental results show that the proposed PFedSIS surpasses previous works to a significant extent.
PFedSIS focuses on the SIS task, which serves as a foundational task supporting numerous downstream applications, such as image-guided automated suturing manipulations [53], tool pose estimation [54], instrument tracking [55], trajectory prediction [56], intraoperative navigation [57], and contributes to the next generation of automation in operating intelligence [58]. Surgical instruments are non-deformable, articulated rigid bodies, which allows SGE to exploit a common shape prior across sites. Segmentation tasks on deformable-shape surfaces, such as tumor segmentation with distinctive abnormal structures will pose a challenge for extracting mutual shape information. We will extend our method to tissue (with distinct shapes) segmentation tasks and use generative methods (e.g. diffusion models), to introduce shape perturbations.
@@ 6 Conclusions
This paper presents PFedSIS, a novel PFL method for SIS, capitalizing on visual trait priors of appearance disparity and instrument shape similarity. It first decouples the embedding heads of MSA and channels of convolution layers into personalized and global via GPD. Then, APE enhances personalized parameters tailored to each site’s appearance representation and utilizes inter-site appearance inconsistency via hypernetwork for personalized parameters update, while SGE enables global parameters to maintain shape information via cross-style consistency and effectively share mutual shape information via shape-similarity update. Experimental results on three publicly available datasets show that the proposed PFedSIS achieves superior performance in both quantitative assessment and visual interpretation, while its computation requirements during prediction remain low and comparable to other methods.